# An Approach for Auto Generation of Labeling Functions for Software Engineering Chatbots

### v1

@@ 3 Approach
Figure 4 presents an overview of our approach, which automates the process of generating LFs. The approach takes queries along with their corresponding intents (labeled data) and queries that need to be labeled (unlabeled data) as inputs. The output of our approach is a set of generated LFs that can be used for labeling user queries. Our approach is composed of three main components: (1) Grouper, tasked with expanding the labeled data by detecting semantic similarities between queries in both the labeled and unlabeled data, (2) Generator, responsible for identifying and extracting patterns from the expanded labeled data to generate LFs, and (3) Pruner, which filters high-quality LFs based on their performance. We detail each component in this section.
Figure 4: Overview of our approach and its components.
3.1 Grouper
Previous studies show that the size and diversity of the labeled dataset directly impact the quality of the generated LFs [26, 54, 55]. This is because a more diverse set of queries leads to the generation of LFs that cover different types of user queries. Similar to prior work, we leverage the unlabeled dataset to expand the labeled dataset [56, 57]. More specifically, the Grouper component groups similar queries in the unlabeled dataset and matches them to an intent, in the labeled dataset based on their semantic similarity. For this purpose, the Grouper component leverages the Sentence-t5-xxl [58] transformer to assess the similarity between queries in both labeled and unlabeled datasets. The Sentence-t5-xxl is an 11-billion parameter open-source sentence-based transformer trained in two stages: first on 2 billion question-answer pairs from Community Q&A sites, then fine-tuned on 275K sentence pairs with human-annotated labels from the Stanford Natural Language Inference dataset [58]. Sentence-t5-xxl has been used in prior work to identify the semantic similarity of text [59, 57]. Using Sentence-t5-xxl, The Grouper component identifies the semantic similarity between each query in the unlabelled dataset (unlabelled query) with each query in the labeled dataset (labeled query).
Next, the Grouper augments the unlabelled query to the intent of the labeled query if the semantic similarity between the two queries is higher than a predefined threshold. Otherwise, the Grouper component abstains from adding the unlabelled query to the labeled dataset. Going forward, we refer to the augmented dataset from Grouper as the expanded dataset. The working example in Figure 4 showcases the Grouper component in grouping query (“Who is the author of config.py?”) from the unlabelled dataset that is similar to the query (i.e., similarity $>$ threshold) in the labeled dataset (“I want to know who created file [map.json]”).
3.2 Generator
The main goal of the Generator is to utilize the expanded dataset from the Grouper component to generate LFs. Prior work shows that NLU platforms better classify intents that contain queries with unique characteristics such as distinct entity types or exclusive words [4]. Therefore, the Generator scans all intents’ queries in the labeled data to extract intents’ characteristics that distinguish them from other intent classes. More specifically, the Generator inspects the following characteristics for each intent in the dataset:
Distinct Entity Type: Some intents contain distinct entity types that are not present in other intents. For example, in Figure 1, the entity type ‘file_name’ appears only in the queries of the FileCreator intent. Some NLUs employ entity types as input for intent classification [60]. To identify the intents with distinct entity types, the Generator component scans all queries within the expanded labeled data, including those labeled by the Grouper component, for entities. Here, we utilize Rasa to detect the entities as it has been shown to perform well for SE tasks [4] and is also open-source. Next, the Generator component computes the entities that are distinct to the various intent classes. Finally, the Generator generates the LFs that can label queries that contain these distinct entity types. In the running example, the Generator generates an LF that labels queries with FileCreator intent if the query contains the ‘file_name’ entity type.
Exclusive words: In addition to distinct entity types, intents that contain exclusive words are easier to identify by NLUs [4]. For example, in Figure 1, the words ‘created’ and ‘creator’ appear only in queries related to the FileCreator intent. To identify exclusive words in the input dataset, the Generator components employ a two-step process. First, it extracts all words from the dataset. This is achieved using a CountVectorizer from the scikit-learn library, which converts the text data into a numerical representation and builds a vocabulary of unique words. The CountVectorizer is configured to extract individual words (unigrams), remove common English stop words, and strip accents for character normalization.
Then, for each word, we compute the ratio of its occurrences within a specific intent class to the total occurrences of that word in the entire dataset. Specifically, the Generator component computes the following ratio:
$Exclusivity(word,intent)=\frac{Occurrences(word,intent)}{TotalOccurrences(word)}$ (1)
where $Occurrences(word,intent)$ is the number of times the word appears in queries associated with a specific intent, and $TotalOccurrences(word)$ is the total number of times the word appears across all intents in the dataset. A ratio closer to 1 indicates that a word is highly exclusive to a particular intent. The ratio becomes our uniqueness threshold, a predefined threshold specified by the user. Once the word exclusivity ratios are computed, the Generator immediately proceeds to generate LFs based on words exceeding the predefined uniqueness threshold. For example in Figure 4, the Generator generates LF 1 that contains the words ‘created’ and ‘author’.
Distinct entity and exclusive words: Intents within a dataset may not always be distinguishable merely by a single distinct entity type or by exclusive words. Instead, intents could be characterized through a specific amalgamation of an entity type and exclusive words. For instance, while the entity type ‘file_name’ and the word ‘developer’ may each appear in various intent classes, their combination could be distinctive for the FileCreator intent, as shown in Figure 1. The Generator examines the dataset for unique combinations where specific entities and words co-occur frequently enough to exceed a user-predefined threshold, indicating a particular intent. Upon identifying such an intent, the Generator creates LFs that classify the query as that intent if it contains both the entity type and exclusive words.
Machine Learning (ML) Generated: Although we employ distinct query characteristics to generate the LFs, there are queries that lack clear distinctive features (e.g., distinct entity type). Thus, to make our approach more generalizable by labeling intents that have different characteristics than the ones in the labeled dataset, we employ an ML approach. Specifically, for each intent, we train five ML classifiers namely, Random Forest, Decision Tree, K-Nearest Neighbors, Logistic Regression, and Support-Vector Machine (SVM) on the expanded dataset. These classifiers have been commonly used within the SE literature [61, 62, 63, 64]. We used the labeled dataset as the training data for the classifiers. The input features for the ML-based LFs are extracted from the labeled queries using the CountVectorizer [65], which converts the text data into a matrix of token counts. Each trained classifier will then serve as an LF.
The output of the Generator is a list of LFs for all intents that have unique characteristics. Figure 5 shows examples of generated LFs. The first LF (lines 2 – 7) inspects whether the query contains the words ‘created’ and ‘author’. If this condition is met, it assigns the label FileCreator to the query intent. Otherwise, it returns ‘ABSTAIN’. The second and third LFs (lines 8 – 20) follow a similar pattern, using the presence of specific entities and words to assign the appropriate intent labels. To enable the users of our approach maintain the generated LFs, we also include the class intent and the type of labeling strategy it employs.
@@ 3.2 Generator
Figure 5: Example JSON representation of LFs.
3.3 Pruner
The Generator produces a range of LFs that vary in quality. Some of the generated LFs might have low accuracy, meaning they frequently mislabel queries when gauged against the intents they intend to label [26]. Additionally, some LFs might overfit to a single data point in the training data. These types of LFs can negatively impact the performance of our approach. Therefore, we devise the Pruner to filter them out from the pool of generated LFs.
To achieve this, the Pruner leverages a portion, called evaluation data, of the expanded dataset (i.e., the output of the Grouper) to evaluate all LFs generated by the Generator. Since the expanded dataset contains labeled queries, the Pruner applies the LFs to the evaluation data and computes the accuracy of each LF. In particular, the Pruner discards LFs with lower accuracy because these LFs are likely to mislabel the queries, which could cause the NLU to be trained on incorrect data, thereby potentially degrading its performance. Another key factor is the coverage, which determines the range of applicability for an LF. Thus, the Pruner discards LFs with low coverage. This action also enhances performance speed by decreasing the number of LFs each query needs to be processed through. Another measure of LF quality is LF support, which refers to the number of queries (i.e., data rows) used to train each LF. LF Support captures the essence of training data diversity and volume used to generate each LF. We hypothesize that a large and diverse training dataset leads to the creation of LFs with higher accuracy and coverage. Coverage and Accuracy are established measures from the Snorkel framework [38]. Moreover. they are used by prior work [26, 55].
Using the calculated characteristics, the Pruner discards LFs that do not meet performance thresholds. It is important to note that the threshold of the accuracy, coverage, and support can be configured by the users of our approach, which provides the flexibility to adapt the approach to different datasets and thus extends its applicability. The final output is a list of high-quality LFs, as shown in Figure 6 ready for use in labeling tasks. In our running example, in Figure 4, LF 1 characteristics are higher than the specified threshold for coverage, accuracy, and support. Therefore, it is retained.
@@ 3.3 Pruner
6 "empirical_accuracy": 0.9,
7 "polarity": 1,
8 "intent_queries": 3
@@
14 "empirical_accuracy": 0.8,
15 "polarity": 1,
16 "intent_queries": 2
@@
Figure 6: Final output of the LFs after the pruner.
4 Case Study Setup
@@
4.1 Datasets
To evaluate the performance of our approach of generating LFs for SE chatbots, we selected four datasets previously used to develop SE chatbots [66, 9, 67, 4]. Table I provides an overview of the number of queries and intents for each of the selected datasets. The selected datasets represent various SE tasks, including seeking information related to software projects and software development tasks. Furthermore, the datasets vary in size, ranging from smaller collections with only a few queries per class to larger sets with hundreds of queries. The details for each dataset with their intents are available in the Appendix. This diversity enables us to assess the effectiveness of our approach across different scales. In particular, we use the following datasets:
TABLE I: Overview of the Selected Datasets Used to Evaluate Our Approach.
@@ 4.1 Datasets
AskGit: AskGit [66] is a chatbot that answers software project-related questions (e.g., “How many commits happened during March 2021?”) on Slack. It is published on GitHub Marketplace so that practitioners can install it on their software projects. AskGit developers brainstormed to create the initial training set for the intents supported by AskGit. Then, they piloted the chatbot with practitioners to gather additional training queries for each intent, expanding their final dataset. This dataset contains 749 queries grouped into 52 intents.
MSA: MSA [9] is a chatbot that assists practitioners in creating microservice architectures by providing answers to questions about microservice environment settings (e.g., “Tell me the server’s environment setting”). The MSA dataset contains 83 queries across eight distinct intents.
Ask Ubuntu: The Ask Ubuntu dataset [67] contains some of the most popular questions from the Ubuntu Q&A community on Stack Exchange. The intents of the collected questions were annotated through Amazon Mechanical Turk. This dataset includes 50 queries (e.g., “What screenshot tools are available?”) divided into four intents (e.g., ‘MakeUpdate’).
Stack Overflow: Ye et al. [68] collected software practitioners’ questions posted under the most popular tags on Stack Overflow. Abdellatif et al. [4] then annotated and categorized these questions (e.g., “Use of session and cookies what is better?”) into different intents. This dataset contains 215 queries grouped into five intents (e.g., ‘LookingForBestPractice’).
4.2 NLU Platform
NLU platforms serve as the backbones for chatbots as they enable chatbots to understand the user’s queries [5, 4, 69]. Typically, chatbot developers resort to off-the-shelf NLUs (e.g., Google Dialogflow) in their chatbot rather than developing an NLU from scratch because it requires both NLP and AI expertise [4]. Among the variety of NLUs, we select the Rasa NLU platform to evaluate the impact of the generated LFs on the NLU’s performance. Our motivation for selecting Rasa is that it is an open-source platform, which makes its internal implementation consistent during our evaluation. Thus, it enables the replicability of our study by other researchers compared to the NLUs on the cloud, whose internal implementations could be changed without any prior notice. Rasa can be installed, configured, and run on local machines, which consumes fewer resources compared to the NLUs that operate on the Cloud. Furthermore, Rasa has been used by prior work to develop SE chatbots [9, 11, 70].
4.3 Evaluation Settings
@@
To determine the optimal values for these thresholds, we conducted a systematic evaluation using the MSRBot dataset [11]. We first evaluated for the Grouper’s threshold, varying from 0.1 to 1.0, while keeping other thresholds constant. We found that a threshold of 0.8 yielded the best overall performance. Next, we evaluated for the Generator’s threshold, varying it from 0.1 to 1.0 in increments of 0.1, while keeping the other thresholds constant. We found that a threshold of 0.8 yielded the best overall performance in terms of labeling accuracy. Finally, we evaluated for the Pruner’s threshold, also varying each from 0.1 to 1.0 in increments of 0.1, while keeping the other thresholds constant. We found that a threshold of 0.7 yielded the best overall performance in terms of labeling accuracy.
@@
The process of applying labels to data, particularly in situations where different LFs produce conflicting labels, is a critical step in our approach. To systematically manage these conflicts, we adopt Snorkel’s Majority Label Voting (MLV) strategy [27]. This choice is informed by manual evaluations and is in alignment with methodologies employed in previous studies [38, 71, 44]. The outcome is labeled data ready for NLU training.
4.4 Performance Evaluation
To evaluate the performance of the generated LFs and NLU’s performance, we also compute the widely used metrics of precision, recall, and F1-score. Precision is the percentage of correctly labeled queries to the total number of labeled queries for that intent (i.e., Precision = $\frac{TP}{TP+FP}$). Recall is the percentage of correctly labeled queries to the total number of queries for that intent in the oracle (i.e., Recall = $\frac{TP}{TP+FN}$). To have an overall performance of the generated LFs, we use the weighted F1-measure that has been used by prior work [72, 73]. More specifically, we aggregate the precision and recall using F1-score (i.e., F1-score = $2\times\frac{Precision\times Recall}{Precision+Recall}$) for each class and aggregate all classes F1-measure using a weighted average, with the class’ support as weights.
Furthermore, we compute the Area Under the ROC Curve (AUC). AUC assesses the model’s ability to distinguish between classes by considering the trade-offs between true and false positive rates [28, 74]. An AUC value above 0.5 suggests that the model is capable of classifying instances more accurately than random guessing, making it a robust metric for assessing performance in diverse classification scenarios. The significance of AUC is emphasized by its application in prior research, particularly in studies dealing with class imbalance [74, 29, 75, 76], including SE studies that involved analyzing imbalanced datasets [77, 78, 79].

### v2

@@ 3. Approach
Figure 4 presents an overview of our approach, which automates the process of generating LFs. The approach takes queries along with their corresponding intents (labeled data) and queries that need to be labeled (unlabeled data) as inputs. The output of our approach is a set of generated LFs that can be used for labeling user queries. Our approach is composed of three main components: (1) Grouper, which expands the initial labeled data by identifying and incorporating semantically similar queries from the unlabeled data pool, (2) Generator, responsible for identifying and extracting patterns from the expanded labeled data to generate LFs, and (3) Pruner, which serves as a quality control filter, evaluating all candidate LFs based on their performance and discarding those that are redundant or ineffective. We detail each component in this section.
Figure 4. Overview of our approach and its components.
3.1. Grouper
Prior work shows that the size and diversity of the initial labeled data directly impact the quality of the generated LFs (Varma and Ré, 2018; Xu et al., 2021; Denham et al., 2022). A larger and more varied set of examples allows for the extraction of more robust and generalizable patterns. Therefore, the first component of our approach, the Grouper, is tasked with expanding the initial labeled dataset. Following established data augmentation practices (Schopf et al., 2020; et al., 2023a), the Grouper enriches the labeled data by incorporating semantically similar queries from the large unlabeled data pool.
To achieve this, the Grouper calculates the semantic similarity between each query in the unlabeled set and every query in the labeled seed set. If the similarity score between an unlabeled query and a labeled one exceeds a predefined threshold, the unlabeled query is assigned the intent of its most similar counterpart and integrated into the labeled data. Queries from the unlabeled set that do not meet this threshold for any intent are not grouped and are set aside for later stages. For assessing semantic similarity, we use the Sentence-t5-xxl model (et al., 2021c), an 11-billion parameter sentence-transformer chosen for its proven effectiveness in text similarity tasks (et al., 2020b, 2023a).
The output of this component is a larger and more diverse labeled dataset, which we refer to as the expanded dataset. This expanded dataset serves as the input for the Generator component. For instance, as illustrated in the example in Figure 4, the Grouper identifies that the unlabeled query “Who is the author of config.py?” is semantically similar to the labeled query “I want to know who created file [map.json]”, and consequently adds it to the corresponding intent class.
3.2. Generator
The Generator is the core component responsible for creating LFs. It takes the expanded dataset from the Grouper and analyzes it to find patterns that distinguish one intent from another. Prior work shows that intents containing unique features, such as distinct words or entity types, are more easily classified by NLU platforms (Abdellatif et al., 2021). Based on this, the Generator employs four distinct strategies to create a comprehensive set of candidate LFs.
To prepare the text for analysis, we first process the queries in the expanded dataset using a CountVectorizer, which builds a vocabulary of unigram features. To prevent data leakage, the vectorizer’s vocabulary is built exclusively from this expanded dataset and is reused for all subsequent feature transformations. The entity types for all queries are extracted using the Rasa NLU platform, which has been shown to perform well on SE tasks (Abdellatif et al., 2021).
3.2.1. Exclusive Word LFs
This strategy identifies words that are highly correlated with a specific intent. For each word in the vocabulary, we calculate an exclusivity score for each intent using Equation 1.
(1) $Exclusivity(word,intent)=\frac{Occurrences(word,intent)}{TotalOccurrences(word)}$
A score close to 1 indicates that a word appears almost exclusively within queries for a single intent. If a word’s exclusivity score for an intent exceeds a predefined threshold, an LF is generated. This LF will vote for that intent if the word is present in a query, and abstain otherwise. For instance, in the example shown in Figure 4, the words ‘created‘ and ‘author‘ are identified as exclusive to the FileCreator intent, resulting in the creation of LF 1.
3.2.2. Distinct Entity LFs
This strategy operates similarly to the word-based one but focuses on entity types. It identifies entities that are unique to a specific intent class within the expanded dataset. For each such distinct entity, an LF is created that votes for the corresponding intent whenever that entity type is detected in a query. For example, as shown in Figure 4, an LF (LF 2) is generated to vote for an intent based on the presence of the ‘file_name‘ entity.
3.2.3. Distinct Word-Entity Combination LFs
This strategy captures more complex patterns where the co-occurrence of a specific word and an entity is a strong signal for an intent, even if the individual components are not exclusive on their own. The Generator searches for these unique combinations across the dataset. If a particular word-entity pair is found to be highly indicative of a single intent, an LF is created that votes for that intent only when both the word and the entity are present. As an example, Figure 4 shows the generation of LF 3 based on the combination of the word ‘author‘ and the ‘file_name‘ entity.
3.2.4. Machine Learning (ML) LFs
While heuristic-based LFs are effective, some intents are best distinguished by more complex, non-linear patterns. To capture these, we also train a suite of standard ML classifiers (Random Forest, Decision Tree, K-Nearest Neighbors, Logistic Regression, and SVM) on the vectorized features of the entire expanded dataset. Each trained classifier is then encapsulated as a single, powerful LF. Unlike the heuristic-based LFs, which typically vote for one specific intent or abstain, these ML-based LFs can classify a query into any of the possible intents, providing broader classification capability.
The combined output of these four strategies is a large set of candidate LFs, as exemplified in Figure 5. This set is then passed to the Pruner for quality control and refinement.
@@ 3.2.4. Machine Learning (ML) LFs
Figure 5. Example JSON representation of LFs.
3.3. Pruner
The Generator produces a large set of candidate LFs, not all of which are equally effective. Some LFs may have low accuracy, others may have very limited applicability (low coverage), and some may overfit to noisy or overly specific patterns in the training data. The Pruner is the final component of our approach, designed to serve as a quality control filter that evaluates and selects only the most robust and generalizable LFs from this initial candidate pool.
The pruning process begins by holding out a portion of the expanded dataset as an evaluation set. Each candidate LF is then applied to this evaluation set to calculate its performance across three key characteristics:
  1. (1)
    Accuracy: To measure an LF’s labeling accuracy, we use the weighted F1-score (Sokolova and Lapalme, 2009). This metric is chosen because it provides a balanced measure of precision and recall, which is crucial for multi-class problems.
  2. (2)
    Coverage: This metric represents the percentage of the evaluation set that an LF labels (i.e., for which it does not abstain). It measures the LF’s range of applicability (et al., 2020a, 2017a).
  3. (3)
    LF Support: We introduce this metric to measure the robustness of an LF’s underlying pattern. It is defined as the number of queries in the expanded dataset that were used to create the LF. A higher support suggests the LF is based on a more common and reliable signal.
Based on these characteristics, the Pruner executes a multi-step filtering algorithm to construct the final set of LFs. First, to ensure that no intent is ignored, the single best-performing LF (based on F1-score) for each intent class is automatically retained. Second, from the remaining pool of LFs, we apply a core quality filter, discarding any LF that has (a) an accuracy below a predefined threshold, (b) a coverage of zero on the evaluation set, or (c) an LF Support of less than two (i.e., was generated from only a single query). This step removes inaccurate LFs, those with no practical applicability, and those highly likely to have overfitted.
The final output is a list of high-quality LFs, as shown in Figure 6, ready for use in labeling tasks. In our running example, in Figure 4, LF 1 characteristics are higher than the specified threshold for coverage, accuracy, and support. Therefore, it is retained.
@@ 3.3. Pruner
6 "accuracy": 0.9,
7 "lf_support": 3
@@
13 "accuracy": 0.8,
14 "lf_support": 2
@@
Figure 6. Final output of the LFs after the pruner.
4. Case Study Setup
@@
4.1. Datasets
To evaluate the performance of our approach of generating LFs for SE chatbots, we selected four datasets previously used to develop SE chatbots (AskGit, [n. d.]; Lin et al., 2020; et al., 2016; Abdellatif et al., 2021). Table 1 provides an overview of the number of queries and intents for each of the selected datasets. The selected datasets represent various SE tasks, including seeking information related to software projects and software development tasks. Furthermore, the datasets vary in size, ranging from smaller collections with only a few queries per class to larger sets with hundreds of queries. The details for each dataset with their intents are available in the Appendix. This diversity enables us to assess the effectiveness of our approach across different scales. In particular, we use the following datasets:
Table 1. Overview of the Selected Datasets Used to Evaluate Our Approach.
@@ 4.1. Datasets
AskGit: AskGit (AskGit, [n. d.]) is a chatbot that answers software project-related questions (e.g., “How many commits happened during March 2021?”) on Slack. It is published on GitHub Marketplace so that practitioners can install it on their software projects. AskGit developers brainstormed to create the initial training set for the intents supported by AskGit. Then, they piloted the chatbot with practitioners to gather additional training queries for each intent, expanding their final dataset. This dataset contains 749 queries grouped into 52 intents.
MSA: MSA (Lin et al., 2020) is a chatbot that assists practitioners in creating microservice architectures by providing answers to questions about microservice environment settings (e.g., “Tell me the server’s environment setting”). The MSA dataset contains 83 queries across eight distinct intents.
Ask Ubuntu: The Ask Ubuntu dataset (et al., 2016) contains some of the most popular questions from the Ubuntu Q&A community on Stack Exchange. The intents of the collected questions were annotated through Amazon Mechanical Turk. This dataset includes 50 queries (e.g., “What screenshot tools are available?”) divided into four intents (e.g., ‘MakeUpdate’).
Stack Overflow: Ye et al. (Ye et al., 2016) collected software practitioners’ questions posted under the most popular tags on Stack Overflow. Abdellatif et al. (Abdellatif et al., 2021) then annotated and categorized these questions (e.g., “Use of session and cookies what is better?”) into different intents. This dataset contains 215 queries grouped into five intents (e.g., ‘LookingForBestPractice’).
4.2. NLU Platform
NLU platforms serve as the backbones for chatbots as they enable chatbots to understand the user’s queries (Prajwal et al., 2019; Abdellatif et al., 2021; Canonico and Russis, 2018). Typically, chatbot developers resort to off-the-shelf NLUs (e.g., Google Dialogflow) in their chatbot rather than developing an NLU from scratch because it requires both NLP and AI expertise (Abdellatif et al., 2021). Among the variety of NLUs, we select the Rasa NLU platform to evaluate the impact of the generated LFs on the NLU’s performance. Our motivation for selecting Rasa is that it is an open-source platform, which makes its internal implementation consistent during our evaluation. Thus, it enables the replicability of our study by other researchers compared to the NLUs on the cloud, whose internal implementations could be changed without any prior notice. Rasa can be installed, configured, and run on local machines, which consumes fewer resources compared to the NLUs that operate on the Cloud. Furthermore, Rasa has been used by prior work to develop SE chatbots (Lin et al., 2020; Abdellatif et al., 2020; Melo et al., 2020).
4.3. Evaluation Settings
@@
To determine the optimal values for these thresholds, we conducted a systematic evaluation using the MSRBot dataset (Abdellatif et al., 2020). We first evaluated the Grouper’s threshold, varying from 0.1 to 1.0, while keeping other thresholds constant. We found that a threshold of 0.8 yielded the best overall performance. Next, we evaluated for the Generator’s threshold, varying it from 0.1 to 1.0 in increments of 0.1, while keeping the other thresholds constant. We found that a threshold of 0.8 yielded the best overall performance in terms of labeling accuracy. Finally, we evaluated for the Pruner’s threshold, also varying each from 0.1 to 1.0 in increments of 0.1, while keeping the other thresholds constant. We found that a threshold of 0.7 yielded the best overall performance in terms of labeling accuracy.
@@
The process of applying labels to data, particularly in situations where different LFs produce conflicting labels, is a critical step in our approach. To systematically manage these conflicts, we adopt Snorkel’s Majority Label Voting (MLV) strategy (et al., 2020a). This choice is informed by manual evaluations and is in alignment with methodologies employed in previous studies (et al., 2017a; Lison et al., 2020, 2021). The outcome is labeled data ready for NLU training.
4.4. Performance Evaluation
To evaluate the performance of the generated LFs and NLU’s performance, we also compute the widely used metrics of precision, recall, and F1-score. Precision is the percentage of correctly labeled queries to the total number of labeled queries for that intent (i.e., Precision = $\frac{TP}{TP+FP}$). Recall is the percentage of correctly labeled queries to the total number of queries for that intent in the oracle (i.e., Recall = $\frac{TP}{TP+FN}$). To have an overall performance of the generated LFs, we use the weighted F1-measure that has been used by prior work (et al., 2019, 2018a). More specifically, we aggregate the precision and recall using F1-score (i.e., F1-score = $2\times\frac{Precision\times Recall}{Precision+Recall}$) for each class and aggregate all classes F1-measure using a weighted average, with the class’ support as weights.
Furthermore, we compute the Area Under the ROC Curve (AUC). AUC assesses the model’s ability to distinguish between classes by considering the trade-offs between true and false positive rates (Yang et al., 2022; Bradley, 1997). An AUC value above 0.5 suggests that the model is capable of classifying instances more accurately than random guessing, making it a robust metric for assessing performance in diverse classification scenarios. The significance of AUC is emphasized by its application in prior research, particularly in studies dealing with class imbalance (Bradley, 1997; Ferri et al., 2009; He and Garcia, 2009; Bekkar et al., 2013), including SE studies that involved analyzing imbalanced datasets (Khatoonabadi et al., 2023a, b; Shatnawi, 2017).